OpenStudy (anonymous):

3x^2+x-2=0 can it be factorized? or must i use qudratic formula?

OpenStudy (anonymous):

Use the formula. It can't hurt.

OpenStudy (anonymous):

i think all quadratic polynomials can be factorized

OpenStudy (anonymous):

I think his question is whether it can be done by "observation" vs. using the formula.

OpenStudy (anonymous):

try \[(x+1) (3 x-2)\]

OpenStudy (anonymous):

if coefficient near x^2 is not 1 or others coefficients is not natural numbers it's hard to factorize by observation

OpenStudy (anonymous):

Yes, it can be factorize.

OpenStudy (asnaseer):

there is a fairly good article here on how to learn to do this: http://mash.dept.shef.ac.uk/Resources/web-factorisingquadratics.pdf

OpenStudy (anonymous):

3(x^2+x/3-2/3)=3[(x+1/6)^2-1/36-2/3)]=3[(x+1/6)^2-25/36] =3(x+1/6-5/6)(x+1/6+5/6)=3(x-2/3)(x+1)=(3x-2)(x+1)..voila